OpenStudy (datman21328):

Particle A starts at A(1,0) and moves in the first upper half of the elipse 4x^2 + y^2=4 at what point is the particle farthest from Starting point A , what is the maximum distance ?

OpenStudy (mww):

This one is a tricky one. I recommend you write distance in terms of a parameter, theta using x = a cos (theta) and y = b sin (theta). Then differentiate distance wrt to theta and find what theta will give you the maximum distance.

OpenStudy (irishboy123):

use the distance formula \(s^2 = (x-1)^2 + y^2\) and replace \(x^2\) with the \(y^2\) using the equation of ellipse so that it becomes an expression in x only then set \(\dfrac{d (s^2)}{dx} = 0\)